OpenStudy (anonymous):

use f(x)=x+2 and g(x)=3x+1 to find specified function (g o f)^-1

OpenStudy (espex):

For this it would appear that you need to compose the inverse of g(f(x))

OpenStudy (anonymous):

(g o f)^-1=g^-1(f(x)) 1. => Find g^-1(x): g(x)=3x+1 So g(g^-1(x))=3*g^-1(x)+1 We know for any function g(g^-1(x))=x .: x=3*g^-1(x)+1 Solving for g^-1 using elementary algebra: g^-1(x)=(x-1)/3 Substituting f(x) into x: g^-1(f(x))=(f(x)+1)/3 ...=(x+2+1)/3 ...=(x/3)+1 .:(g o f)^-1=(x/3)+1